# Finite element approximation of the stationary Navier-Stokes problem with non-smooth data

### v1

@@ 6 Conclusions
The finite numerical approximation of the stationary Navier-Stokes equations with non-smooth data have been analyzed. We have compared the original problem with its corresponding version using smooth data, closely related to the original data, and the associated finite element approach. Convergence and optimal a priori error estimates have been rigorously proven under suitable assumptions regarding the domain and finite element spaces.
The reported numerical experiments for the lid-driven cavity flow problem, with different finite element methods, have illustrated the effectiveness of our approach.

### v2

@@ 6 Conclusions
The finite numerical approximation of the stationary Navier-Stokes equations with non-smooth data have been analyzed. We have compared the original problem with its corresponding version using smooth data, closely related to the original data, and the associated finite element approach. Convergence and optimal a priori error estimates have been rigorously proven under suitable assumptions regarding the domain and finite element spaces. The reported numerical experiments for the lid-driven cavity flow problem, with different finite element methods, have illustrated the effectiveness of our approach.
[…]
Acknowledgements We thank Ricardo G. Durán for his helpful references on a priori estimates for the Stokes problem, and Ariel L. Lombardi for his valuable suggestions regarding numerical implementation.
Declarations
Funding. This work was partially supported by grant PICT 2018-3017, Agencia Nacional de Promoción de la Investigación, el Desarrollo Tecnológico y la Innovación (Argentina), DOI http://dx.doi.org/10.13039/501100021778 and by grant 20020170100056BA, Universidad de Buenos Aires, DOI
http://dx.doi.org/10.13039/501100005363.
Competing interests. The authors have no competing interests to declare that are relevant to the content of this article.
Data availability. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.